I have attempted this problem about 6 times now, and I always get stuck :( I'm supposed to solve for a but I cant figure it out! Can someone help me? Also, if you could include steps that would be wonderful :)
solve for a
a - (3a-6)/6 = (3a+2)/3

Answer by rfer(16322)

3(a-(3a-6)/6)=3a+2
3(a-(3a-1)=3a+2
3(a-3a+1)=3a+2
3a-9a+3=3a+2
-9a=-1
a=1/9